DEIMOS Status Report
Science Steering Committee
April 23, 2001
By Sandy Faber
This is a briefer-than-normal report since Dave Cowley is not here right now and I am acting as Project Manager, which is keeping me busy.
FLEXURE/IMAGE MOTION/GRATING SYSTEM:
1) The stiffening of the grating box seems to have worked, based on dial gauge measurements of the structure when rotated. Flexure of the grating box per se is about 10 times less than previously. That would mean an image motion of about 3 px pk-pk from this source alone.
2) We have been working through the grating drive/mounting mechanisms systematically, with the following results:
- Slider 2 (mirror) meets all tests. It clamps repeatably in all PAs, and image rotation under PA change is predicted to be <1 px pk-pk at the edge of the detector. (This rotation was measured relative to the grating box. There may be other sources of image rotation. The spec on total image rotation is 1 px pk-pk at the edge of the detector. Image rotation is not correctable by the FC system.)
- Slider 3 is heavier (carries a grating). We worked through each step of the mounting/clamping process and made several improvements. The challenge was to achieve complete isolation from the drive system when clamped, as the drive system flexes and potentially puts stresses into the sliders. This now seems to be successful. As of yesterday we had the following results:
* When clamped, the total rotation of the dummy Al blank inside the slider implies a spectral rotation of <1 px pk-pk at edge of detector. As above, there may be other sources of image rotation, so this is a partial number.
* When held by the drive screw before clamping, the slider sags at the opposite end by +/-0.025-in in PA 0 and 180. This nearly exceeds the capture range of the mounting pin inside the hole in the grating box. Inserting this pin is the first step in grating clamping. However, the offset is repeatable and could be corrected by applying a small PA-dependent correction to the nominal slider motor-insertion location. Alternatively, we could install an optical fiducial on the pin end of the slider. Either way, the problem does not seem serious.
* Slider 3 clamps successfully in PA 0, 180, and 270 once the pin has found the hole. All gaps close and the slider is well seated. The slider does NOT clamp successfully yet in PA 90, there being a residual gap of a few thousandths of an inch at the stop that controls rotation. This problem is undiagnosed as yet.
* The slide drive system has been carrying the weight of sliders 2,4, and 4 during the above tests. The additional weight of slider 5 when installed is not a large perturbation on the present system and should not cause a problem.
- Slider 4 needs extensive bracing at the part that holds it to the drive screw, with the result that sag at the pin hole is extensive. It looks like there is room to install this bracing, though the fit will be tight. Such sag if uncorrected would affect the ability of slider 4 to find the pin hole but should not affect the flexure or image rotation once properly clamped. We are looking at bracing designs now.
- We have officially deleted the large slider 1 with the 8x12 in grating capability. Its replacement, slider 5, will look rather like slider 3, and its design should be simple once the kinks have been worked out of slider 3.
- The principal engineer of the grating system is Jack Osborne, who is losing time owing to a severe illness in his family. Vern Wallace has taken over the debugging of the system, which is delaying his work on the carriage mover (trolley) system).
- Our goal is to have sliders 2 and 3 fully debugged and ready for the next round of optical testing by May 1.
3) We have not done much optical testing since the last report. However, a small set of data was taken with the Cohu camera by shining light out through the camera and back off various mirrors mounted on the camera and the grating box. The main finding is that we have a loose optical element in the camera, which should cause a total image motion of about 12 px pk-pk in direct-viewing mode. We think this is the largest single remaining source of flexure. The suspected elements are either Element 3 or Multiplet 4. It has not been decided when we will disassemble the camera to fix this.
4) Stiffening of the detector stage, carried out since our last report, reduced the flexure within the dewar by about a factor of two. We had hoped for a factor of 3-4. The total remaining motion in each coordinate is about 4 px pk-pk.
5) Summary of total remaining sources of image translation (does not include rotation): Some are guesses and will be much better measured during the next round of optical tests. It is not known how the phases add.
- Moving optical element in camera: * 12 px pk-pk
- Collimator (based on early tests last fall): ? 4 px pk-pk
- Detector stage in dewar: 4 px pk-pk
- Grating box + slider (suspected, not measured): 8 px pk-pk
The * indicates a source of flexure we intend to fix. The ? indicates a source of flexure we may be able to improve. The remaining two sources are probably as good as we can make them in the foreseeable future.
- Net flexure estimate (worst case, assuming phases add) 16
px pk-pkz
The total capture range of the FC system is 20 px pk-pk. We are still distressingly
close to this. We'd like to get total flexure down to 10 px pk-pk to be comfortably
within the capture range.
6) Image quality:
- The previous round of optical tests in Dec-Jan revealed significant coma in the radial direction. The suspected cause was incorrect spacing between Multiplet 4 and the field flattener. Increasing this by 0.060-in would be needed to cure the effect, according to Zemax.
- We later found that we left a 0.030-in shim out, accounting for half the effect.
- The other half is plausibly attributed to comatic aberrations at room temperature, which should vanish when the camera operates at 0 C.
- We checked this conclusion with a large series of Cohu camera images taken at different positions in the FOV at different shim thicknesses. Increasing the shim thickness by 0.030-in beyond normal indeed produced the best images. These looked pretty good, although the Cohu has such limited dynamic range that we cannot calculate enclosed energies reliably.
- We'll know more about image quality in the next round of optical tests, where we plan to adjust the shim thickness. This is easy to do, and a final value will be adopted after optical testing at the operating temperature at Keck.
7) Dewar/detector/controller:
- We converted successfully to 16-amp mode during February. However, the B amplifier on CCD#3 did not function properly. This was unexpected, as it had functioned well in earlier lab tests (in a different dewar). This suggests that our CCD installation procedures may still damage CCDs, despite all precautions. We will convene a tiger team to review our procedures once more before installing the red detector.
- Otherwise, the blue mosaic is operating well, and the dark current and ion pump glow are acceptably low.
- An intermittent flaw was discovered on the Motorola Power PC board that is CARA's recommended replacement for the Force board in the VME crate. It appears only at the high video rates we experience. Motorola sent a spare, which had the same problem. The real cause may in fact lie in a Chrislin 256 Mb memory board, which issues an intermittent, erroneous error signal at high data rates that paralyzes the Motorola. Lack of a working Motorola board increases the video data transfer time from 40 sec to 70 sec, lengthening the effective mosaic RO time also to 70 sec. Solutions are being explored.
8) Red science mosaic:
- One more excellent Lot 14 device was received from MIT/LL, bringing the number of superb devices to 7. Eight are needed for the red mosaic. Two reasonable spares are also in hand.
- We are going to ask Gerry Luppino to give us the remaining red CCD.
9) Slitmask system:
- At our Jan report, slitmasks would not retract fully at PA 90. This has been fixed by testing a more powerful spring plunger. However, we can no longer use the shiniest aluminum for slitmasks, as it is soft and easily deformed by the plunger. This will reduce the amount of light going into the TV guider from the slitmask region, but the amount of this loss has not been calculated. The stationary offset guider mirror is unaffected.
- The slitmask system will be removed to install the new plungers and make final adjustments next week.
10) PA rotation:
- We now have keyword control of all aspects of rotation, including raw position and velocity.
- The reference encoder for PA rotation angle is the Renashaw encoder on the rear bearing, which has two readout heads. The two heads agree at all PA's to within 20 arcsec. The nominal spec for PA accuracy is 17 arcsec, which we meet easily if the two readouts are averaged, and nearly meet even with one head.
- The raw encoder readings will need at least sin theta and sin 2*theta correction terms for decenter and out-of-roundness of the Renashaw tape. It is our intention to find the rough magnitude of these terms using precision levels in the lab and find final coefficients on the sky curing commissioning.
- Al Honey of CARA is coming this week to help in getting the PA system to talk to the CARA DCS.
11) Carriage mover (trolley):
- Our present plan is to lay about 20 feet of tracks in the shop and test the trolley system fairly realistically. Most of the mechanical components are in hand, and much of the design has been done.
- Work is being delayed on this system at the moment because the engineer, Vern Wallace, has been transferred to grating-system testing.
12) Remaining optical elements: Several of these are behind schedule and may have to be delivered late in the fall.
- We received a disappointing quote from Cleveland Crystal for the Solgel coating of the front window: $31,000. The original quote had been $21,000 from Coherent, which we rejected because it seemed too high at the time. We will now contact Livermore immediately to see if they can do it more cheaply, and recontact Coherent to see if they are still interested.
- Dave Hilyard is going to fabricate our broadband filters, at no charge to DEIMOS. Bill Brown will apply Mg2F coatings in the Lick coating chamber. I am designing the final filter glasses and passbands now. We are behind schedule on this activity. We have added a z-band filter because of the high sensitivity of the Lot 14 devices at 1 micron.
- A major outstanding question is how to coat the order-blocking filters and the broadband clear filter. The wavelength ranges of these elements, delta lambda/lambda, are large, from 2 to 2.6 and are beyond the capabilities of the our facility.
- At one time, Coherent quoted $20,000 to coat the clear filter. We could conceivably save quite a lot of money by going to two clear filters, each with smaller delta lambda, but this would necessitate more filter handling and less observing flexibility. Livermore Solgel is again a possiblity.
13) Software:
- The final version of the Lickserve2 CCD RO software is nearly finished, including final keyword control of the PANE (subraster) keywords and temperature control system keywords.
- We have keyword control at the Galil level of the scissors and insert subsystems of the slitmask system. The first test of the merged system is planned for today.
- The piezo and barcode software was written and lab tested. It is awaiting debugging on the real instrument.
- The one remaining system without much motor control software is the grating system, though pieces of it have been worked on. Software progress has been delayed by the nearly constant mechanical work on this system.
- The data-taking computer Celeste was successfully reconfigured to run Solaris 2.8. The same upgrade on the instrument computer DEITRADER failed because KROOT had been incorrectly configured at CARA. This cost us nearly two weeks to track down but has now been fixed.
- Our main motor control programmer, Dean Tucker, has been afflicted by severe back pain and now by a separate problem that requires surgery. He will be out for at least a week. The grating control software is on the critical path, and we expect significant delays. His absence could be covered by Kibrick working with Deich and Allen if necessary, but at the cost of further lost time.
- The other remaining chunk of software is the flexure compensation system software, which has not been started. The final system will clearly not be ready at the nominal Preship Review date in late June. We have complied a draft list of minimum performance requirements as of that date for negotiation with CARA.
14) Shipping:
- Our shipper is Matson Lines. We toured their facilities in Oakland and boarded a ship to view firsthand the transfer process from ship to shore. They seem like good people and to have our problem well in hand.
- Two vessels will be needed for the main structure: a ship from Oakland to Honolulu, and a barge from Honolulu to Kawaihae. The structure will be carried below decks to Hawaii but above decks on the barge. It will also sit for a couple of days on the pier in Honolulu, where a protectivecover will again be needed.
- An extensive list of performance requirements was assembled for the shipping crate for the main structure, including weather protection, internal humidity, and shock loading. Engineer Jerry Cabak is designing the shipping container.
- The camera will be disassembled into five pieces for shipping to Hawaii. This is necessary to encase each optical multiplet snugly in layers of foam. The present centration of the bodies is recorded using dowel pins and should be easy to recover during assembly.
15) CARA interface:
- The weight of DEIMOS is at issue. CARA was counting on no more than 18,000 lb; we estimate more like 20,000 lb. The exact weight is uncertain by 10% and will be certified by using calibrated load cells, which have been purchased. About 500 lb. have been removed by moving counterweights and replacing steel hatches with aluminum. CARA's current position is that they will be able to accommodate a 20,000 lb instrument. They have undertaken extensive bracing of the Nasmyth to accomplish this. We very much appreciate their flexibility on this point. Note added after the SSC meeting: CARA representatives said that they had NOT committed as yet to accommodating 20,000 lb. so this item is still open.
- CARA visited in February and wrote an extensive report of items that they would like to be improved. Most of these revolve around convenience and safety. A draft response is being negotiated. We are doing our best to respond to nearly all items.
16) Schedule:
- As of last week, we were technically about two weeks behind the last SSCschedule on January 25. However, we will probably slip still more due to the illness of Dean Tucker (motor control software for the grating system) and the extended absence of Jack Osborne (and consequent delay on the trolley system). My estimate (not cleared with Dave Cowley) is that the Preship Review will be delayed about a month, to the end of July.
- Date of Readiness Review with CARA. This is a Preship Review checkout that should take place about 6 weeks before the Preship Review itself. The current date is the end of May. Realistically, we probably won't make this. More like end of June?
- Ways to save time:
* Postpone construction of slider 5 until after the Preship Review and install it in Hawaii: a good idea, since this slider is expected to be straightforward.
* Postpone final fixes on slider 4 to Hawaii: not a good idea, since the slide drive connection to 4 is tricky, and it is much easier to debug this here. It needs to be connected to the whole system to make sure it works.
* Postpone fixing the slipping optical element in the camera until camera disassembly for shipping: needs further evaluation.
* Postpone checkout of the trolley system: possibly a good idea. Needs further discussion. I don't have a feeling for the risks.
* Postpone coating of the broadest filters and/or the front window until fall: unavoidable.
* Postpone final implementation of the FC system: unavoidable if either a June or July Preship schedule is to be met. The minimum elements of the system at Preship then need to be defined. We are working on this.